Known DT online problems

LINK to the List of Known problems in 2009


FAQ: DSS: As seen in the PVSS (LV & HV)

As seen in the PVSS (LV & HV)

  • 1) YB-2 S7 MB4 SL3 L1 off it stays in ramping status (yellow) switching HV to standby.

As seen in the DCS_GUI

  1. YB2 S4 MB2 FE max temperature is ~100 degrees FAQ
  2. YB2 S2 MB2 ROB 3 TDC 1 fatal (L1 buffer parity error). (== FED774/ROS2/ROB 9 tdc fatal in DQM) - Not critical but if you can send a robreset whenever the run is halted (and before the DAQ reconfigures again) it would be good.
  3. YB2 S1 MB2 going through Secondary Port. Ok, donīt worry.


As seen in the ROS status pages

Total number of ROB's blocked: 0 (YB+2 S1 MB3 is not blocked with present configuration)


As seen in the DDU status pages

See DDU status pages


As seen in the TSC


As seen in the TS DTTF

For DTTF masks check:

https://twiki.cern.ch/twiki/bin/view/CMS/DTTriggerRuns2010

  • 1) YB-2 S4 MB4 masked.

As seen in the DQM online


Occupancy problems


Occupancy problems due to RO problems

Aproximately 0 channels lost.


Occupancy problems due to HV problems

See occupancy plots from run 119017 and 130269

Aproximately, 178 chamber channels lost.

  • 1) YB+1 S4 MB3 SL3 L1 ch 0 to 32 and 50 to 57 missing ---(23-Mar-09) TJB#08, L1 s HVB ABCD 32 channels
  • 2) YB0 S1 MB1 SL2 --- (28-Sep-09) TJB#9 L2 S HVB G 10 channels. Looks like the entire SL is inefficient, not very visible.
  • 3) YB0 S4 MB1 SL1 L4 ch 0 to 15 --- (20-Apr-09) TJB#10, L4 s HVB AB 16 channels
  • 4) YB0 S4 MB4 SL1 L4 ch 0 to 15 less entries, probably noise --- (23-Jul-09) L4 MB4/03 1/4 layer OFF 16 channels. Strip problem in MB4/3.
  • 5) YB0 S9 MB4 SL1 L3 all --- (12-Jun-09) TJB#13, L3 s HVB EFGH 32 channels. Strip problem.
  • 6) NEW! YB0 S11 MB1 SL3: - L1 ch 41-49 and L3 ch 0-16 --- e-log ID 173102 Entry time: Tue February 23 12:46:32 2010
  • 7) YB-1 S4 MB2 SL2 L1 ch 33 to 48 missing ---- (10-Jul-09) TJB#02 L1 s HVB EF 16 channels. Strip problem.
  • 8) YB-1 S10 MB4-10(9) SL1 L3 ch 17 to 24 missing ---(24-Aug-09) TJB#11 10/9L3w0 HVB C 8 channels


Occupancy problems due to "unknown" problems (likely FE problems)

See occupancy plots from run 119017, 130269 and 131423

Aproximately, 52 chamber channels lost (plus 228 if YB+2 S5 MB1 SL2 is not working)

  1. Occupancy problems YB-1 S1 MB3 SL1 L1 ch29 to 32 (TP and cosmics) masked due to FEB very noisy. It is masked usually, if it is not (INFORM TO MASK IT).
  2. YB-1 S12 MB4 SL1 all layers ch 81 to 84 (TP and in cosmics getting worst) Run 109861 already there. In Run 132482 is OK.
  3. YB+2 S2 MB4 SL1 L2 and L4 ch 0 to 16 (missing in TP, cosmics fine) Run 99563 already there. Cable TP line 1 disconnected. Problem gone in Run 138869 (June 2010).
  4. NEW! YB-2 S1 MB2 SL1 all layers. Channels from 53 to 56 empty. (TP and cosmics). ok in 2009
  5. YB+2 S5 MB1 SL2 (TP and cosmics) Entire Superlayer . Ok in 2009. RECOVERED since run 133908, since after technical stop 26 april. Back empty again beggining of may, then recovered at some point in june. It comes and goes.
  6. NEW! YB+1 S7 MB1 SL3 all layers ch 0 to 3 (TP and cosmics) Run 131423 already there. Probably FEB dead, curious high noise in SL1 L2 and L3.


Timeboxes problems


Noise problems

  1. Noise problems YB-1 S1 MB3 SL1 L1 ch29 to 32 (TP and cosmics) masked due to FEB very noisy. It is masked usually, if it is not (INFORM TO MASK IT).
  2. YB0 S6 MB2 SL2 L2 ch 10 and L1 ch 9 Seen in March 2010
  3. YB0 S6 MB1 SL1 L3 ch 49 noisy cells
  4. YB1 S3 MB3 SL1 L4 ch 30 noise at the end of the timebox
  5. YB2 S8 MB1 SL2 L2 ch 57 noisy cell
  6. YB3 S12 MB1 SL1 L3 ch 2 noisy cell


Trigger problems

For DTTF masks check:

https://twiki.cern.ch/twiki/bin/view/CMS/DTTriggerRuns2010

  • 1) YB+1 S9 MB2: Lack of HH triggers (DCC only)
  • 2) YB0 S10+S09 MB2: Inefficiency (DCC+DDU)
  • 3) No DCC info YB-2 S04 MB4


As seen in the DT prompt offline

  • 1) Intermitent TDCFatal error in YB-2 S2 MB2 (ROB3), seen as ROB9 of ROS2 in Prompt offline analysis results, for instance Run 142130


Miscellaneous

  • YB-1 S6 MB1: a TRB is OFF (TRB6, theta trigger). Generates a "Vdd too low error" on PVSS (17.8 A). Errors on DCS GUI have been masked.


-- GianniMasetti - 05-Nov-2009

-- IgnacioRedondo - 08-Aug-2010

Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r4 - 2010-08-19 - IgnacioRedondo
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Sandbox All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &Đ 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback